OpenStudy (anonymous):

WILL FAN AND MEDAL The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 132 = one-halfx + 2(x - 9) What is the regular price of a child's ticket? $60 $57 $51 $42

OpenStudy (cj49):

1/2(x+4(x-9)=132 1/2(x+4x-36)=132 x+4x-36=264 5x=300 x=60 60-9=51

OpenStudy (anonymous):

Thank you! I have one more! Ill put it in another post!

OpenStudy (cj49):

np